OpenStudy (anonymous):

What divided by 74 would equal -3?

OpenStudy (phi):

call the unknown number "x" x divided by 74 = -3 or \[ \frac{x}{74}= -3 \] if you multiply both sides of the equation by 74 you get \[ 74 \cdot \frac{x}{74} = -3 \cdot 74 \] can you simplify that ?

OpenStudy (anonymous):

-222?